OpenStudy (anonymous):

Find an equivalent equation in rectangular coordinates 1) r= 5 + 8 sin theta 2) r= 6/ (3+ cos theta)

OpenStudy (tkhunny):

Perform substitutions and simplify. There are only three.

OpenStudy (anonymous):

i tried but its hard...can u show me one?

OpenStudy (yttrium):

\[r = 5 + 8 \sin \theta\] \[x^2 + y^2 = 5 +\frac{ y }{ x^2 + y^2 }\] Just simplify this one

OpenStudy (anonymous):

w8

OpenStudy (yttrium):

I forgot to write 8. it should be 8y in the denominator

OpenStudy (anonymous):

i thought sin theta = y/r

OpenStudy (anonymous):

8 sin theta = 8y/r = 8y / square root ( x^2 + y^2)

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

oh right!!

OpenStudy (yttrium):

sorry. same as in the left hand side.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

r = square root ( x ^2 + y^2)

OpenStudy (yttrium):

yea. so generally. the answer is x^2 + y^2 = 5 sqrt(x^2+y^2)+ 8y

OpenStudy (anonymous):

just keep the square root?

OpenStudy (yttrium):

no